# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix D Welfare formulas used in Section 3
$\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-,\iota}&=\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\frac{d}{2}\log\left(1+\frac{(1-R)(\beta+\widehat{y}^{2}R)}{\beta R(1+2\widehat{y})}\right)\\ &\qquad+\frac{1}{2}\frac{\beta((\widehat{y}+1)^{2}+\beta)^{2}}{(\widehat{y}+1)^{4}\left(\beta R(1+2\widehat{y})+(1-R)\left(\beta+\widehat{y}^{2}R\right)\right)}\left\|P_{X}^{-1/2}\left(\psi_{I,0}-\psi_{U,0}\right)\right\|^{2}\end{split}$
and
@@
$\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-,\iota}&=\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\log\left(1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\right)\\ &\qquad+\frac{1}{2}\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{(\widehat{y}+1)^{4}(\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}+\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2}))}\bigg{\|}P_{X}^{-1/2}\left(\psi_{I,0}-\psi_{U,0}\right)\bigg{\|}^{2}.\end{split}$
Proof of Proposition D.5.
We start with the insider. From Lemma D.2 and (87) the interim certainty equivalent becomes
$\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})&=\psi_{I,0}^{\prime}\mu_{X}+\widehat{y}\sqrt{\frac{R(1-R)}{\beta}}\psi_{I,0}^{\prime}P_{X}^{-1/2}\bigg{(}\mathcal{E}_{N}-\sqrt{\frac{\beta}{R(1-R)}}\frac{\beta+(\widehat{y}+1)^{2}}{\lambda_{U}\widehat{y}(\widehat{y}+1)^{2}}P_{X}^{-1/2}\widehat{\Pi}\\ &\qquad-\sqrt{\frac{\beta}{R(1-R)}}\frac{\beta}{(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{)}+\frac{(1-R)\widehat{y}^{2}}{\beta(1+\widehat{y})}\bigg{\|}\mathcal{E}_{N}-\sqrt{\frac{\beta}{R}}\frac{1}{\widehat{y}}\mathcal{E}_{G}\\ &\qquad-\sqrt{\frac{\beta}{R(1-R)}}\frac{\beta+(\widehat{y}+1)^{2}}{\lambda_{U}\widehat{y}(1+\widehat{y})^{2}}P_{X}^{-1/2}\widehat{\Pi}-\sqrt{\frac{\beta}{R(1-R)}}\frac{\beta+R(\widehat{y}+1)^{2}}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}\bigg{\|}^{2}.\end{split}$
We now use Lemma D.3, but unlike in Proposition D.4, we set $\mathcal{E}_{A}=\mathcal{E}_{N}$ and $\mathcal{E}_{B}=\mathcal{E}_{G}$. This gives
$\begin{split}C_{1}&=\psi_{I,0}^{\prime}\mu_{X};\quad C_{2}=0;\quad C_{3}=\frac{(1-R)\widehat{y}^{2}}{\beta(1+2\widehat{y})};\quad C_{4}=\sqrt{\frac{\beta}{R}}\frac{1}{\widehat{y}},\\ V_{1}&=\sqrt{\frac{R(1-R)}{\beta}}\widehat{y}P_{X}^{-1/2}\psi_{I,0},;\quad V_{2}=\sqrt{\frac{\beta}{R(1-R)}}\left(\frac{\beta+(\widehat{y}+1)^{2}}{\lambda_{U}\widehat{y}(1+\widehat{y})^{2}}P_{X}^{-1/2}\widehat{\Pi}+\frac{\beta}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}\right),\\ V_{3}&=-\sqrt{\frac{\beta}{R(1-R)}}\left(\frac{\beta+(\widehat{y}+1)^{2}}{\lambda_{U}\widehat{y}(1+\widehat{y})^{2}}P_{X}^{-1/2}\widehat{\Pi}+\frac{\beta+R(1+\widehat{y})^{2}}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}\right).\end{split}$
As $C_{2}=0$ there are a number of cancellations and hence
$\begin{split}\gamma_{U}\textrm{CE}^{I}_{0-,\iota}&=C_{1}-V_{1}^{\prime}V_{2}+\frac{d}{2}\log\left(1+C_{3}+C_{3}C_{4}^{2}\right)+\frac{1}{2}\frac{C_{3}}{1+C_{3}+C_{3}C_{4}^{2}}\|V_{1}-V_{3}\|^{2}-\frac{1}{2}\|V_{1}\|^{2}.\end{split}$
As in the price taking case, the remainder of the proof simplifies the expression for the certainty equivalent. We already know $C_{1}$. Next
$V_{1}^{\prime}V_{2}=\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}\psi_{I,0}^{\prime}P_{X}^{-1}\widehat{\Pi}+\frac{\beta\widehat{y}}{(1+\widehat{y})^{2}}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}.$
Continuing
$\begin{split}1+C_{3}+C_{3}C_{4}^{2}&=1+\frac{(1-R)(\beta+R\widehat{y}^{2})}{\beta R(1+2\widehat{y})}=\frac{\mathbf{V}}{\beta R(1+2\widehat{y})},\end{split}$
where
$\mathbf{V}:=\beta R(1+2\widehat{y})+(1-R)(\beta+R\widehat{y}^{2}).$
Next,
$\begin{split}\|V_{1}\|^{2}&=\frac{R(1-R)\widehat{y}^{2}}{\beta}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0},\\ \frac{C_{3}}{1+C_{3}+C_{3}C_{4}^{2}}\|V_{1}-V_{3}\|^{2}&=\frac{\widehat{y}^{2}}{\beta(1+\widehat{y})^{4}\mathbf{V}}\bigg{\|}\left(\beta^{2}+R(1+\widehat{y})^{2}(\beta+(1-R)\widehat{y})\right)P_{X}^{-1/2}\psi_{I,0}\\ &\qquad+\frac{\beta(\beta+(1+\widehat{y})^{2})}{\lambda_{U}\widehat{y}}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
At this point, we may conclude
$\begin{split}\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-,\iota}&=\psi_{I,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\log\left(1+\frac{(1-R)(\beta+R\widehat{y}^{2})}{\beta R(1+2\widehat{y})}\right)\\ &\qquad-\left(\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}\psi_{I,0}^{\prime}P_{X}^{-1}\widehat{\Pi}+\frac{\beta\widehat{y}}{(1+\widehat{y})^{2}}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}\right)-\frac{1}{2}\frac{R(1-R)\widehat{y}^{2}}{\beta}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}\\ &\qquad+\frac{\widehat{y}^{2}}{2\beta(1+\widehat{y})^{4}\mathbf{V}}\bigg{\|}\left(\beta^{2}+R(1+\widehat{y})^{2}(\beta+(1-R)\widehat{y})\right)P_{X}^{-1/2}\psi_{I,0}\\ &\qquad+\frac{\beta(\beta+(1+\widehat{y})^{2})}{\lambda_{U}\widehat{y}}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}+\frac{1}{2}\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}.\end{split}$
The $\widehat{\Pi}^{\prime}P_{X}^{-1}\widehat{\Pi}$ terms are
$\frac{1}{2}\frac{\beta(\beta+(1+\widehat{y})^{2})^{2}}{\lambda_{U}^{2}\mathbf{V}(1+\widehat{y})^{4}}.$
The $\psi_{I,0}^{\prime}P_{X}^{-1}\widehat{\Pi}$ terms are
$\begin{split}&-\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}+\frac{\widehat{y}(\beta+(1+\widehat{y})^{2})\left(\beta^{2}+R(1+\widehat{y})^{2}(\beta+\widehat{y}(1-R))\right)}{\lambda_{U}(1+\widehat{y})^{4}\mathbf{V}},\\ &\qquad=\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{4}\mathbf{V}}\left(-\mathbf{V}(1+\widehat{y})^{2}+\beta^{2}\widehat{y}+R\widehat{y}(1+\widehat{y})^{2}(\beta+\widehat{y}(1-R))\right).\end{split}$
The terms within the parentheses on the right evaluate to
$\beta^{2}\widehat{y}-\beta(1+R\widehat{y})(1+\widehat{y})^{2}=-\frac{\beta}{\lambda_{U}}(\beta+(1+\widehat{y})^{2}),$
where the equality follows from (69) and $\lambda_{U}+\lambda_{I}=1$. Therefore, the $\psi_{I,0}^{\prime}P_{X}^{-1}\widehat{\Pi}$ terms are
$-\frac{\beta(\beta+(1+\widehat{y})^{2})^{2}}{\lambda_{U}^{2}\mathbf{V}(1+\widehat{y})^{4}}.$
Lastly, the $\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}$ terms not on the first line are
$\begin{split}&\frac{1}{2}-\frac{\beta\widehat{y}}{(1+\widehat{y})^{2}}-\frac{R(1-R)\widehat{y}^{2}}{2\beta}+\frac{\widehat{y}^{2}\left(\beta^{2}+R(1+\widehat{y})^{2}(\beta+\widehat{y}(1-R))\right)^{2}}{2\beta\mathbf{V}(1+\widehat{y})^{4}},\\ &\qquad=\frac{1}{2}-\frac{\widehat{y}(2\beta^{2}+R(1-R)\widehat{y}(1+\widehat{y})^{2})}{2\beta(1+\widehat{y})^{2}}+\frac{\widehat{y}^{2}\left(\beta^{2}+R(1+\widehat{y})^{2}(\beta+\widehat{y}(1-R))\right)^{2}}{2\beta\mathbf{V}(1+\widehat{y})^{4}}.\end{split}$
We create the common denominator $2\beta\mathbf{V}(1+\widehat{y})^{4}$. The numerator is
$\begin{split}&(1+\widehat{y})^{2}\mathbf{V}\left(\beta(\widehat{y}+1)^{2}-\widehat{y}\left(2\beta^{2}+\widehat{y}R(1-R)(\widehat{y}+1)^{2}\right)\right)\\ &\qquad+\widehat{y}^{2}\left(\beta^{2}+R(\widehat{y}+1)^{2}(\beta+\widehat{y}(1-R))\right)^{2}.\end{split}$
Let us group by powers of $(\widehat{y}+1)$. The fourth order terms are
$\begin{split}&\widehat{y}^{2}R^{2}(\beta+\widehat{y}(1-R))^{2}+\mathbf{V}(\beta-\widehat{y}^{2}R(1-R))\\ &\qquad=\widehat{y}^{2}R^{2}(\beta+\widehat{y}(1-R))^{2}+\left(\beta+2\beta R\widehat{y}+(1-R)R\widehat{y}^{2}\right)(\beta-\widehat{y}^{2}R(1-R))\\ &\qquad=\beta^{2}\left(1+R\widehat{y}\right)^{2}.\end{split}$
The second order terms are
$\begin{split}&-2\beta^{2}\widehat{y}\mathbf{V}+2\beta^{2}R\widehat{y}^{2}(\beta+\widehat{y}(1-R))=-2\beta^{3}\widehat{y}(1+R\widehat{y}).\end{split}$
The $0^{th}$ order term is $\beta^{4}\widehat{y}^{2}$. The numerator is thus
@@
$\beta^{2}\left((1+\widehat{y})^{2}(1+R\widehat{y})-\beta\widehat{y}\right)^{2}=\frac{\beta^{2}(\beta+(1+\widehat{y})^{2})^{2}}{\lambda_{U}^{2}},$
where the equality follows from (69). Therefore, the $\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}$ terms not on the first line are
$\frac{\beta(\beta+(1+\widehat{y})^{2}}{2\lambda_{U}^{2}(1+\widehat{y})^{4}\mathbf{V}}.$
The result follows for the insider $I$ in view of (89). We next consider $U$. From Lemma D.2 and (87) the interim certainty equivalent becomes
$\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0,\iota}&=\psi_{U,0}^{\prime}\mu_{X}+\sqrt{\frac{(1-R)(\beta+R(1+\widehat{y})^{2})}{\beta}}\frac{\widehat{y}}{1+2\widehat{y}}\psi_{U,0}^{\prime}P_{X}^{-1/2}\bigg{(}\mathcal{E}_{H_{\iota}}\\ &\qquad+\sqrt{\frac{\beta}{(1-R)(\beta+R(1+\widehat{y})^{2})}}\frac{\lambda_{I}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}\\ &\qquad-\sqrt{\frac{\beta}{(1-R)(\beta+R(1+\widehat{y})^{2})}}\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}\widehat{y}(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\qquad+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{2\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\bigg{\|}\mathcal{E}_{H_{\iota}}+\sqrt{\frac{\beta(\beta+R(1+\widehat{y})^{2}))}{1-R}}\frac{\widehat{y}}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}\\ &\qquad-\sqrt{\frac{\beta(\beta+R(1+\widehat{y})^{2}))}{1-R}}\frac{1}{\lambda_{I}(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
We now use Lemma D.3 with $\mathcal{E}_{A}=\mathcal{E}_{H_{\iota}}$ as well as
$\begin{split}C_{1}&=\psi_{U,0}^{\prime}\mu_{X};\quad C_{2}=0;\quad C_{3}=\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}};\quad C_{4}=0,\\ V_{1}&=\sqrt{\frac{(1-R)(\beta+R(1+\widehat{y})^{2})}{\beta}}\frac{\widehat{y}}{1+2\widehat{y}}P_{X}^{-1/2}\psi_{U,0},\\ V_{2}&=-\sqrt{\frac{\beta}{(1-R)(\beta+R(1+\widehat{y})^{2})}}\left(\frac{\lambda_{I}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}-\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}\widehat{y}(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\right),\\ V_{3}&=\sqrt{\frac{\beta(\beta+R(1+\widehat{y})^{2}))}{1-R}}\left(\frac{\widehat{y}}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}-\frac{1}{\lambda_{I}(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\right).\end{split}$
As $C_{2}=C_{4}=0$ there are a number of cancellations and hence
@@
$\begin{split}\gamma_{U}\textrm{CE}^{I}_{0-,\iota}&=C_{1}-V_{1}^{\prime}V_{2}+\frac{d}{2}\log\left(1+C_{3}\right)+\frac{1}{2}\frac{C_{3}}{1+C_{3}}\|V_{1}-V_{3}\|^{2}-\frac{1}{2}\|V_{1}\|^{2}.\end{split}$
As with $I$, the remainder of the proof simplifies the expression for the certainty equivalent. We already know $C_{1}$. Next
$V_{1}^{\prime}V_{2}=-\frac{\lambda_{I}\widehat{y}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})^{2}}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{I,0}+\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})}\psi_{U,0}^{\prime}P_{X}^{-1}\widehat{\Pi}.$
Continuing
$\begin{split}1+C_{3}&=1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}=\frac{\mathbf{V}}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}},\end{split}$
where
(90) $\mathbf{V}:=\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}+\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2}).$
Next,
$\begin{split}\|V_{1}\|^{2}&=\frac{(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{\beta(1+2\widehat{y})^{2}}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0},\\ \frac{C_{3}}{1+C_{3}}\|V_{1}-V_{3}\|^{2}&=\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}}\\ &\qquad\times\bigg{\|}P_{X}^{-1/2}\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}\psi_{U,0}-\frac{\widehat{y}}{(1+\widehat{y})^{2}}\psi_{I,0}+\frac{1}{\lambda_{I}(1+\widehat{y})}\widehat{\Pi}\right)\bigg{\|}^{2}.\end{split}$
At this point, we may conclude
$\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-,\iota}&=\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\log\left(1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\right)\\ &\qquad+\left(\frac{\lambda_{I}\widehat{y}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})^{2}}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{I,0}-\frac{\beta+(1+\widehat{y})^{2}}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})}\psi_{U,0}^{\prime}P_{X}^{-1}\widehat{\Pi}\right)\\ &\qquad+\left(\frac{1}{2}-\frac{(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\beta(1+2\widehat{y})^{2}}\right)\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}\\ &\qquad+\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}\\ &\qquad\times\bigg{\|}P_{X}^{-1/2}\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}\psi_{U,0}-\frac{\widehat{y}}{(1+\widehat{y})^{2}}\psi_{I,0}+\frac{1}{\lambda_{I}(1+\widehat{y})}\widehat{\Pi}\right)\bigg{\|}^{2}.\end{split}$
The remainder of the proof shows this value coincides with that in the statement of the Lemma. Here, unlike with $I$ we cannot leave $\widehat{\Pi}$ as is until the very end. Instead we substitute in for $\widehat{\Pi}$ as in (89). Doing this, and simplifying, we obtain
$\begin{split}\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-,\iota}&=\psi_{U,0}^{\prime}\mu_{X}-\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\log\left(1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\right)\\ &\qquad-\frac{\lambda_{I}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})^{2}}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{I,0}-\frac{\beta+(1+\widehat{y})^{2}}{(1+2\widehat{y})(1+\widehat{y})}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}\\ &\qquad+\left(\frac{1}{2}-\frac{(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\beta(1+2\widehat{y})^{2}}\right)\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}\\ &\qquad+\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}\\ &\qquad\times\bigg{\|}P_{X}^{-1/2}\left(\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}+\frac{\lambda_{U}}{\lambda_{I})(1+\widehat{y})}\right)\psi_{U,0}+\frac{1}{(1+\widehat{y})^{2}}\psi_{I,0}\right)\bigg{\|}^{2},\end{split}$
The $\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{I,0}$ terms are
$\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}(1+\widehat{y})^{4}},$
which is consistent with that in the statement of the lemma. As for the $\psi_{I,0}^{\prime}P_{X}^{-1}\psi_{U,0}$ terms, we claim
@@
(91) $\begin{split}&-\frac{\lambda_{I}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})^{2}}+\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}(1+\widehat{y})^{2}}\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}+\frac{\lambda_{U}}{\lambda_{I}(1+\widehat{y})}\right)\\ &\qquad=-\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}(1+\widehat{y})^{4}},\end{split}$
which is consistent with the Lemma statement. Indeed, cancelling common terms and re-arranging, (91) is the same as
@@
$\begin{split}&\frac{1}{\lambda_{U}(1+2\widehat{y})}=\frac{\beta\lambda_{I}(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}}\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}+\frac{\lambda_{U}}{\lambda_{I}(1+\widehat{y})}+\frac{1}{(1+\widehat{y})^{2}}\right),\end{split}$
or equivalently
$\begin{split}&\lambda_{U}(\beta+R(1+\widehat{y})^{2})\bigg{(}\lambda_{I}(1-R)\widehat{y}(1+\widehat{y})^{2}+\lambda_{U}\beta(1+2\widehat{y})(1+\widehat{y})+\beta\lambda_{I}(1+2\widehat{y})\bigg{)}\\ &\qquad=(1+\widehat{y})^{2}\mathbf{V}\\ &\qquad=\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}(1+\widehat{y})^{2}+\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})(1+\widehat{y})^{2}\\ &\qquad=\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}(1+\widehat{y})^{2}+\lambda_{U}\lambda_{I}(1-R)(R(1+\widehat{y})^{2}\widehat{y}-\beta(1+\widehat{y}))(1+\widehat{y})^{2},\end{split}$
where the last equality uses (83). There is an $\lambda_{U}\lambda_{I}R(1-R)\widehat{y}(1+\widehat{y})^{4}$ term common to both sides. Eliminating this term, the remaining terms each have a $\beta$ and $\lambda_{U}$ in them, which we can also eliminate. This leaves
$\begin{split}&(\beta+R(1+\widehat{y})^{2})\bigg{(}\lambda_{U}(1+2\widehat{y})(1+\widehat{y})+\lambda_{I}(1+2\widehat{y})\bigg{)}+\lambda_{I}(1-R)\widehat{y}(1+\widehat{y})^{2}\\ &\qquad=\lambda_{U}R(1+2\widehat{y})^{2}(1+\widehat{y})^{2}-\lambda_{I}(1-R)(1+\widehat{y})^{3}.\end{split}$
Note that $(1+\widehat{y})^{3}+\widehat{y}(1+\widehat{y})^{2}=(1+\widehat{y})^{2}(1+2\widehat{y})$. This allows us to eliminate one copy of $(1+2\widehat{y})$ from both sides, to obtain
$\begin{split}&(\beta+R(1+\widehat{y})^{2})\bigg{(}\lambda_{U}(1+\widehat{y})+\lambda_{I}\bigg{)}+\lambda_{I}(1-R)(1+\widehat{y})^{2}\\ &\qquad=\lambda_{U}R(1+2\widehat{y})(1+\widehat{y})^{2}=\lambda_{U}R\left((1+\widehat{y})^{3}+y(1+\widehat{y})^{2}\right)\end{split}$
There is a common $\lambda_{U}R(1+\widehat{y})^{3}$ term. Eliminating this term gives
$\begin{split}&\lambda_{I}(\beta+R(1+\widehat{y})^{2})+\lambda_{U}\beta(1+\widehat{y})+\lambda_{I}(1-R)(1+\widehat{y})^{2}=\lambda_{U}Ry(1+\widehat{y})^{2}.\end{split}$
The left side above simplifies to
$\beta+\lambda_{U}\beta\widehat{y}+\lambda_{I}(1+\widehat{y})^{2}.$
Using (83) again, the right side above simplifies to
$\lambda_{I}((1+\widehat{y})^{2}+\beta)+\lambda_{U}\beta(1+\widehat{y})=\beta+\lambda_{U}\beta\widehat{y}+\lambda_{I}(1+\widehat{y})^{2},$
and hence (91) holds. Lastly, for the $\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}$ terms, we claim

### v2

@@ B.4. Utility formulas used in Section 3
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-,\iota}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{R}{2(1+2\widehat{y})}\left|\frac{1-R}{R}P_{X}^{1/2}G_{0}-\widehat{y}P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2}}\right].$
If we re-write this using the notation of (77), (81) we obtain
@@
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-,\iota}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{R}{2(1+2\widehat{y})}\left|\frac{1-R}{R}\left(\mathcal{E}_{X}+\sqrt{\frac{R}{1-R}}\mathcal{E}_{I}\right)-\widehat{y}\sqrt{\frac{1-R}{\beta R}}\mathcal{E}_{N}\right|^{2}}\right].$
We can write
$\frac{1-R}{R}\left(\mathcal{E}_{X}+\sqrt{\frac{R}{1-R}}\mathcal{E}_{I}\right)-\widehat{y}\sqrt{\frac{1-R}{\beta R}}\mathcal{E}_{N}=\sqrt{\frac{(1-R)(\beta+\widehat{y}^{2}R)}{\beta R^{2}}}\mathcal{E},$
where $\mathcal{E}\sim N(0,1_{d})$ under ${\mathbb{Q}}_{0}$ $N(0,1_{d})$. This implies
@@
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-,\iota}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\left(1+\frac{(1-R)(\beta/R+\widehat{y}^{2})}{(1+2\widehat{y})\beta}\right)^{-1/2},$
which gives the formula in (90). We now consider the uninformed trader. From (42), (43), (47) and (54) we obtain
@@
$e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0,\iota}}=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}p_{\iota}-\widehat{\psi}_{U,\iota}^{\prime}(X-p_{\iota})}\big{|}\sigma(H_{\iota})\right]=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(H_{\iota})\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U,\iota}^{\prime}(X_{0}-p_{0,\iota})}\big{|}\sigma(H_{0,\iota})\right].$
From (75) we find
$\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U,\iota}^{\prime}(X_{0}-p_{0,\iota})}\big{|}\sigma(H_{0,\iota})\right]=e^{-\frac{1}{2}\left|(P_{U,\iota}+P_{X})^{-1/2}\left(P_{U,\iota}H_{0,\iota}-(P_{U,\iota}+P_{X})p_{0,\iota}(H_{0,\iota})\right)\right|^{2}}.$
From (60), (62) and (79) calculation shows
@@
$\begin{split}&(P_{U,\iota}+P_{X})^{-1/2}\left(P_{U,\iota}H_{0,\iota}-(P_{U,\iota}+P_{X})p_{0,\iota}(H_{0,\iota})\right)\\ &\qquad=\sqrt{\frac{(1-R)^{2}(1+\widehat{y})^{2}(\beta-R\widehat{y}(1+\widehat{y}))^{2}}{(\beta+R(1+\widehat{y})^{2})R(\beta+(1+\widehat{y})^{2})(1+2\widehat{y})^{2}}}P_{X}^{1/2}H_{0,\iota}.\end{split}$
From (78) we deduce
$(1+\widehat{y})^{2}(\beta-R\widehat{y}(1+\widehat{y}))^{2}=\left(\frac{\lambda_{I}}{\lambda_{U}}(\beta+(1+\widehat{y})^{2})\right)^{2},$
which gives
@@
$\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U,\iota}^{\prime}(X_{0}-p_{0,\iota})}\big{|}\sigma(H_{0,\iota})\right]=e^{-\frac{1}{2}\frac{\lambda_{I}^{2}(1-R)^{2}(\beta+(1+\widehat{y})^{2})}{\lambda_{U}^{2}R(\beta+R(1+\widehat{y})^{2})(1+2\widehat{y})^{2}}\left|P_{X}^{1/2}H_{0,\iota}\right|^{2}}.$
The interim certainty equivalent for $U$ in (89) follows from (87) and (47). For the ex-ante certainty equivalent we must compute